# Signature Kernel Scoring Rule as Spatio-Temporal Diagnostic for Probabilistic Forecasting

### v1

@@ 4 Training a Multi-step Probabilistic Forecaster
We employ generative models to provide probabilistic forecasts. These models, specifically generative neural networks, work by transforming randomly sampled latent variables into samples on the required output space. They achieve this by learning a mapping $G$ from a low dimensional simple distribution, $z\sim\mathcal{N}(0,I)$, to our observations $x$, $G:z\rightarrow x$. The application of deep neural networks allows our generative model to learn and capture the complex distribution by applying non-linear functions and working with hierarchical representations. This type of model is highly appropriate to the problem of weather forecasting as generative networks predict probabilistic outputs by design by generating samples as opposed to point predictions. However, the common way to train a generative network, via generative adversarial training, is unstable, suffers from mode collapse, and has a training objective which has not been extended to temporal data (Salimans et al., 2016). As a result, we implement the predictive-sequential (prequential) scoring rule to train generative neural networks to minimise scoring rules, first introduced by Lorenzo Pacchiardi et al. (Pacchiardi et al., 2024). We first consider the one step ahead prequential SR formulation, before extending to a prequential SR for pathlengths $l$ for sliding window generative models.
@@ 4.1 Generative Models and the Prequential Scoring Rule
The general forecasting problem is that we are given a set of $k$ past realisations, and we want to predict the next observation, $k+1$. In general we can consider any specific future observation $k+l$, but for simplicity, we consider the immediately sequential step. Having observed $t$ previous values, we employ a generative model, with parameters $\phi$ to give us a forecast distribution conditioned on the previous window of k observations: $P^{\phi}_{t+1}(\cdot|y_{t-k+1:t})$. We can then evaluate the performance of the forecast at time step $t+1$ with a given score S:
@@
Indeed, for such a one step forecast, optimisation over $\phi$ through minimisation of the above term selects for the parameters that yield forecasts whose average score across the training data is optimal. Under conditions of stationarity and mixing for $(\textbf{Y}_{1},\textbf{Y}_{2},\ldots,\textbf{Y}_{T})$, it has been proven that the empirical minimiser $\hat{\phi}$ is consistent, i.e. converges to the minimiser of the expected prequential SR (Pacchiardi et al., 2024). Therefore, to optimise $\phi$ via stochastic gradient descent we need to obtain unbiased estimates of $\nabla(P^{\phi}_{t+1}(\cdot|y_{t-k+1:t}),y_{t+1})$, which occur when S is strictly proper and defined by expectation over $P^{\phi}$. Critically, this is true of the signature kernel scoring rule, seen in terms of expectation above in Equation 4, as is standard for a general kernel score.
@@
The full sliding window generation process is displayed in Figure 4. Recall that each prediction $\hat{y}_{l}$ is an ensemble, consisting of $e$ ensemble members, where each member requires a latent variable $z$ for each time step. In fact, for learning purposes, the matrix of latent variables $\mathbf{Z}$ is static, resulting in a unique noise term for each combination of ensemble member $e$ and path position $l$, regardless of initialisation time $t$. The result of generation is $e$ paths of length $l$, which can finally be evaluating with a chosen score.
@@
The objective function for the patched signature kernel score, for patches $p$ $\in$ $\mathcal{P}$, initialisation time $t$, path length $l$, ensemble size $m$ is as follows:
@@
Unlike in evaluation, many standard weather models, such as the ones in Table 1, do not train on latitude weighted scores. As a result, we do not need to take latitude slices and can consider global paths or patches. Patches also disrupt spatial structure, but include both latitude and longitudinal correlations and may capture local behaviour more effectively. Additionally, lower dimensionality is helpful to reduce numerical stability in this method, addressed further in Appendix C. Both approaches were tested, and patches proved slightly more effective. We consider 24 patches of size 16x16, with three latitude and eight longitude initialisation positions, accounting for longitude wrapping. Each patch has a dimension size of 256, eight times smaller than the global 2048 dimensions. Patches have considerable overlap to ensure smooth prediction behaviour and cover the equivalent of three globe areas. Overlap naturally occurs in the non-polar regions, which implicitly latitude weights to a moderate degree.
4.2 Model architecture
We employ a U-NET architecture for the generative network. U-NET consists of encoders, which are a series of convolution layers followed by max pooling. Each level increases the number of feature maps allowing the model to learn complex features. A decoder follows, which upsamples the features along with concatenating the results of previous encoders using skip connections. This allows the decoder to reuse features from earlier layers, preserving both low-frequency and high-frequency information. Additionally, we consider a bottleneck layer between the encoder and decoder where the latent variables are introduced, to make this process generative. We consider a U-NET architecture with depths of 32, 64, 128, and 256 respectively, illustrated in 5. We implement the same window size of 10 previous time steps, and apply the same sliding window prequential generation. Therefore, the U-NET takes a time input of size 10, and gives an output for size 1, across the 64 by 32 spatial coordinates.
Figure 5: Diagram of the implemented U-NET model architecture. The input size is equivalent to the window size of 10. The output size is 1. The U-NET has depths of 32, 64, 128, and 256. The latent variables, introduced between the encoder and decoder, are denoted by Z.
@@ 4.2 Model architecture
We consider geopotential 500 hPa on the roughest resolution of 64 by 32 longitude-latitude regions. We use global data from 2010 to 2018, with a 6:2:1 train, validation, test split using the years 2010-2015, 2016-2017, and 2018 respectively. Optimal training would involve a significantly larger number of years to capture longer term weather patterns. However, these data choices were made to balance learning ability and computational time. We consider all experiments with a maximum of 100 epochs and validation early stopping in increments of 5 epochs past 20. We train across 4 path lengths (2, 5, 10, 15) and each across 9 learning rates, logarithmically spaced between 1e-2 and 1e-6. All use the RBF static kernel with $\sigma=1$, dyadic order of 1, and initial context window size of 10. Maximum epoch time was never reached during training. The model chosen for evaluation is based on the lowest final validation loss reached. We consider a batch size of 16 patches due to memory constraints on the GPU. Experimental results with a one-step prequential scoring rule on WeatherBench data showed equivalent metrics of calibration error, NRMSE, and $R^{2}$ across training ensemble numbers of three to fifty (Pacchiardi et al., 2024). Therefore, we train with three ensemble members to minimise computation time with no compromise to performance. All metrics are evaluated on an evaluation ensemble size of 200.
4.3 Training Results
@@
Table 2 displays the latitude-weighted metric results of training on WeatherBench Data across four different path lengths. As expected, smaller path lengths results in stronger performance, particularly notable in NRMSE and $R^{2}$. At path length 15, the model’s $R^{2}$ is roughly equal to zero, suggesting equal performance to mean climatology. Nonetheless, the results across path lengths are strong, with low NCRPS and low calibration error. For all path lengths, NCRPS is substantially less than 1, implying a forecast error smaller than the historical spread of climatology. Additionally, calibration error appears to plateau around 0.2 across path lengths 5 to 15.
WeatherBench Training Results
@@
Table 2: Evaluation metrics across path length for signature kernel trained models.
The following figures illustrate the performance of the signature kernel trained models across path length. Figures 6 and 7 showcase ensemble spread for one realisation, with random variations between members to capture the forecast distribution. Most discrepancies exist in the temperate zone where there is distinct boundary behaviour with polar regions. Remaining figures showcase one ensemble member across respective path lengths of 5, 10, and 15. Errors are minimal for each initial forecast and build with increasing path length as expected.
Figure 6: Ensemble predictions for one initialisation generated from the path length 2 model.
Figure 7: Ensemble differences for one initialisation generated from the path length 2 model.
Figure 8: Path comparison for one ensemble member generated from the path length 5 model.
Figure 9: Path differences for one ensemble member generated from the path length 5 model.
Figure 10: Path comparison for one ensemble member generated from the path length 10 model.
Figure 11: Path differences for one ensemble member generated from the path length 10 model.
Figure 12: Path comparison for one ensemble member generated from the path length 15 model.
Figure 13: Path differences for one ensemble member generated from the path length 15 model.
@@ 5 Conclusion
We highlight the following extensions to this work: first, greater investigation into handling numerical stability through scaling and static kernels will allow for more robust training results across learning rates and extending to higher dimensionality weather forecasts. Secondly, our current approach of probabilistic non-adversarial training is not common, and implementation and adaptation into current weather forecasting frameworks is necessary for an appropriate analysis of performance.
@@ A.1 RMSE
with $\theta_{i}^{u}$ and $\theta_{i}^{l}$ being the upper and lower latitude bounds for the grid cell with latitude index (center) $i$. Polar regions, which would require a different formula, are excluded from the 64 by 32 5.625 resolution ERA5 dataset which is used for training.
For use in evaluation to standardise RMSE results across variables, we introduce the normalised root mean square error NRMSE [Pacchiardi et al., 2024]:
@@ A.2 $R^{2}$
where $\bar{o}^{v}$ represents the observation mean. $R^{2}$ in this formulation is bounded above by 1, when forecasts equal observations across all times, $f^{v}_{t,i,j}=o^{v}_{t,i,j}$, but is unbounded from below, due to the absence of an intercept in the neural network. $R^{2}=0$ suggests a deterministic model which does no better than predicting the mean for every observation, with $R^{2}<0$ therefore suggesting poor model fit. We include this metric in evaluation to compliment RMSE in assessing accuracy of the forecast mean predictions.
@@ A.3 CRPS
We next consider standard probabilistic scoring rules, designed for ensemble forecasts. The continuous ranked probability score (CRPS), which is strictly proper, is particularly relevant for the probabilistic forecasting of rainfall, due to rainfall having a positive point mass at 0. While other scoring rules for continuous variables are restricted to predictive densities, CRPS instead uses predictive cumulative distributions [Gneiting and Raftery, 2007]. For deterministic forecasts, the CRPS reduces to MAE. The standard formulation for CRPS for a variable $v$, with the cumulative distribution function of its probabilistic forecast F is as follows:
@@
We evaluate the NCRPS across specific positions, and normalise by the standard deviation of observations. This normalisation is used to assess the quality of the forecast across scale of the data, along with adding greater interpretation to CRPS results. A NCRPS greater than 1 represents a forecast error larger than the typical variation, standard deviation, of the observations. NCRPS smaller than 1 means the forecast error is smaller than the typical variation, suggesting reasonable forecasting performance, especially as NCRPS decreases to zero. This relates the performance to the historical spread of climatology in the region, which will be roughly equal to 1.

### v2

@@ 4 Training a Multi-step Probabilistic Forecaster
We employ generative models to provide probabilistic forecasts. These models, specifically generative neural networks, work by transforming randomly sampled latent variables into samples on the required output space. They achieve this by learning a mapping $G$ from a low dimensional simple distribution, $z\sim\mathcal{N}(0,I)$, to our observations $x$, $G:z\rightarrow x$. The application of deep neural networks allows our generative model to learn and capture the complex distribution by applying non-linear functions and working with hierarchical representations. This type of model is highly appropriate to the problem of weather forecasting as generative networks predict probabilistic outputs by design, by generating samples as opposed to point predictions. However, the common way to train a generative network, via generative adversarial training, is unstable, suffers from mode collapse, and has a training objective which has not been extended to temporal data (Salimans et al., 2016). As a result, we implement the predictive-sequential (prequential) scoring rule to train generative neural networks to minimise scoring rules, first introduced by Lorenzo Pacchiardi et al. (Pacchiardi et al., 2024). We first consider the one step ahead prequential SR formulation, before extending to a prequential SR for pathlengths $l$ for sliding window generative models.
@@ 4.1 Generative Models and the Prequential Scoring Rule
The general forecasting problem is that we are given a set of $k$ past realisations, and we want to predict the next observation, $k+1$. In general, we can consider any specific future observation $k+l$, but for simplicity, we consider the immediately sequential step. Having observed $t$ previous values, we employ a generative model, with parameters $\phi$ to give us a forecast distribution conditioned on the previous window of k observations: $P^{\phi}_{t+1}(\cdot|y_{t-k+1:t})$. We can then evaluate the performance of the forecast at time step $t+1$ with a given score S:
@@
Indeed, for such a one step forecast, optimisation over $\phi$ through minimisation of the above term selects for the parameters that yield forecasts whose average score across the training data is optimal. Under conditions of stationarity and mixing for $(\textbf{Y}_{1},\textbf{Y}_{2},\ldots,\textbf{Y}_{T})$, it has been proven that the empirical minimiser $\hat{\phi}$ is consistent, i.e. converges to the minimiser of the expected prequential SR (Pacchiardi et al., 2024). Therefore, to optimise $\phi$ via stochastic gradient descent, we need to obtain unbiased estimates of $\nabla(P^{\phi}_{t+1}(\cdot|y_{t-k+1:t}),y_{t+1})$, which occur when S is strictly proper and defined by expectation over $P^{\phi}$. Critically, this is true of the signature kernel scoring rule, seen in terms of expectation above in Equation 1, as is standard for a general kernel score.
@@
The full sliding window generation process is displayed in Figure 3. Recall that each prediction $\hat{y}_{l}$ is an ensemble, consisting of $e$ ensemble members, where each member requires a latent variable $z$ for each time step. In fact, for learning purposes, the matrix of latent variables $\mathbf{Z}$ is static, resulting in a unique noise term for each combination of ensemble member $e$ and path position $l$, regardless of initialisation time $t$. The result of generation is $e$ paths of length $l$, which can finally be evaluated with a chosen score.
Observed SequenceGenerative Net $h_{\phi}]$Observation WindowLatent Variables $\mathbf{Z}$$y_{t-k+1}$$y_{t-k+2}{\scriptstyle\cdots}$$y_{t}$$y_{t+1}$$y_{t+2}{\scriptstyle\cdots}$$y_{t+l}$$\hat{y}^{1}_{t+1}$$\hat{y}^{1}_{t+2}{\scriptstyle\cdots}$$\hat{y}^{1}_{t+l}$$\hat{y}^{2}_{t+1}$$\hat{y}^{2}_{t+2}{\scriptstyle\cdots}$$\hat{y}^{2}_{t+l}{\scriptstyle\cdots}{\scriptstyle\cdots}{\scriptstyle\cdots}{\scriptstyle\cdots}$$\hat{y}^{e}_{t+1}$$\hat{y}^{e}_{t+2}{\scriptstyle\cdots}$$\hat{y}^{e}_{t+l}$$\mathcal{N}(0,1)$$h_{\phi}]$Draw $z^{1}_{1},z^{1}_{2},\ldots z^{1}_{e}$$(y_{t-k+2}:y_{t},\,\hat{y}_{t+1})$$\mathcal{N}(0,1)$$h_{\phi}]$Draw $z^{2}_{1},z^{2}_{2},\ldots z^{2}_{e}$$(y_{t},\,\hat{y}_{t+1}:\hat{y}_{t+l-1})$$\mathcal{N}(0,1)$$h_{\phi}]$Draw $z^{l}_{1},z^{l}_{2},\ldots z^{l}_{e}$$z^{1}_{1}$$z^{1}_{2}{\scriptstyle\cdots}$$z^{1}_{l}$$z^{2}_{1}$$z^{2}_{2}{\scriptstyle\cdots}$$z^{2}_{l}{\scriptstyle\cdots}{\scriptstyle\cdots}{\scriptstyle\cdots}{\scriptstyle\cdots}$$z^{e}_{1}$$z^{e}_{2}{\scriptstyle\cdots}$$z^{e}_{l}$
@@
The objective function for the patched signature kernel score, for patches $p$ $\in$ $\mathcal{P}$, initialisation time $t$, path length $l$, and ensemble size $m$, is as follows:
@@
4.2 Minimum Prequential Signature Kernel Score Multistep forecaster for WeatherBench 2
@@
Unlike in evaluation, some standard weather models, such as the ones in Table A.1, do not train on latitude weighted scores (Pathak et al., 2022)(Nguyen et al., 2023). As a result, we are not forced to take latitude slices and can consider global paths or patches. Patches also disrupt spatial structure, but include both latitude and longitudinal correlations and may capture local behaviour more effectively. Additionally, lower dimensionality is helpful to reduce numerical stability in this method, addressed further in Section 4.3. Both approaches were tested, and patches proved slightly more effective. We consider 24 patches of size 16x16, with three latitude and eight longitude initialisation positions, accounting for longitude wrapping. Each patch has a dimension size of 256, eight times smaller than the global 2048 dimensions. Patches have considerable overlap to ensure smooth prediction behaviour and cover the equivalent of three globe areas. Overlap naturally occurs in the non-polar regions, which implicitly latitude weights to a moderate degree.
@@
We employ a U-NET architecture for the generative network, details of which is provided in the Appendix I. For training, we consider geopotential 500 hPa on the roughest resolution of 64 by 32 longitude-latitude regions. We use global data from 2010 to 2018, with a 6:2:1 train, validation, test split using the years 2010-2015, 2016-2017, and 2018, respectively. Optimal training would involve a significantly larger number of years to capture longer term weather patterns. However, these data choices were made to balance learning ability and computational time. We consider all experiments with a maximum of 100 epochs and validation early stopping in increments of 5 epochs past 20. We train across 4 path lengths (2, 5, 10, 15) and each across 9 learning rates, logarithmically spaced between 1e-2 and 1e-6. All use the RBF static kernel with $\sigma=1$, dyadic order of 1, and initial context window size of 10. Maximum epoch time was never reached during training. The model chosen for evaluation is based on the lowest final validation loss reached. We consider a batch size of 16 patches due to memory constraints on the GPU. Experimental results with a one-step prequential scoring rule on WeatherBench data showed equivalent metrics of calibration error, NRMSE, and $R^{2}$ across training ensemble numbers of three to fifty (Pacchiardi et al., 2024). Therefore, we train with three ensemble members to minimise computation time with no compromise to performance. All metrics are evaluated on an evaluation ensemble size of 200.
@@
Table 1 displays the latitude-weighted metric results of training on WeatherBench Data across four different path lengths. As expected, smaller path lengths result in stronger performance, particularly notable in NRMSE and $R^{2}$. At path length 15, the model’s $R^{2}$ is roughly equal to zero, suggesting equal performance to mean climatology. Nonetheless, the results across path lengths are strong, with low NCRPS and low calibration error. For all path lengths, NCRPS is substantially less than 1, implying a forecast error smaller than the historical spread of climatology. Additionally, calibration error appears to plateau around 0.2 across path lengths 5 to 15.
Training Results for WeatherBench 2 (Geopotential 500 hPa)
@@
Table 1: Evaluation metrics across path length for minimum prequential signature kernel score model for Geopotential 500 hPa from 2010 to 2018.
In appendix I, we provide more figures illustrating the performance of the signature kernel trained models across path length. The discrepancies between ensemble members mostly exist in the temperate zone, where there is distinct boundary behaviour with polar regions, and errors are minimal for each initial forecast and build with increasing path length as expected.
4.3 Numerical Instability
The largest issue with training implementation is the instability of computation, with different scales and static kernels causing issues with Inf/Nan values or scores converging to 1. In multiple test instances, the signature kernel score appears to be calculating correctly for small prediction lags, but then displays abnormal converging behaviour. Theoretical and experimental results were limited in application to this context. The generalised data augmentation method proposed by James Morril et al. (Morrill et al., 2021) demonstrated results of different approaches varying significantly by dimensionality of the path, with eight being the largest dimension tested. Moreover, results were only conducted for path length $l\gg d$, the dimension of the data. In this implementation, the dimensionality of the path is 2048 in evaluation and 256 in training.
This type of path is not covered by existing literature. Nonetheless, extending these results proved successful. Standard normalising using the mean and standard deviation of only the observation data is conducted to allow comparisons between models with equal scaling. A further scaling constant of $\frac{1}{\sqrt{I\cdot J}}$ is applied, which is one over the square root of the longitude and latitude dimensions. Despite no theoretical backing, and with pre-processing rescaling being primarily considered for the truncated signature instead (Morrill et al., 2021), this addition solved detectable instability issues for ensemble models and across increasing resolution, verified up to the standard 240 by 121 resolution. The RBF kernel is considerably more robust to numerical instability issues in both evaluation and training than other kernels tested. The linear identity kernel was initially desired for handling high dimensional time series without complex feature mapping, to simplify the interpretation of the kernel, but numerical instability issues remain in training. We believe this instability arises from two compounding factors: the choice of static kernel directly controlling the magnitude of the loss landscape that occurs when predicted paths deviate significantly from observations, and the Goursat PDE solver accumulating numerical error when path differences are large (Salvi et al., 2021). This issue is evident with the linear and polynomial kernel, see Appendix G, compared to the RBF, which was more robust via its exponential decay moderating the loss landscape. As a result, we note that for reasonable kernel choices, these instability issues are not present in the primary use of the signature kernel score as a diagnostic, where predicted paths remain closely tied to the distribution of observed weather.
@@ 5 Conclusion
We highlight the following extensions to this work: first, greater investigation into handling numerical stability through scaling and static kernels will allow for more robust training results across learning rates and extending to higher dimensionality weather forecasts. Secondly, our current approach of probabilistic non-adversarial training is not common, and implementation and adaptation into current weather forecasting frameworks are necessary for an appropriate analysis of performance. A further discussion of extensions and limitations of this work is provided in Appendix K.
@@ Appendix A Evaluation Metrics
Table A.1 displays a summary analysis of the scoring rules in use among eight of the leading ML data-driven weather forecasting models between 2022-2024, determined by inclusion in the Google WeatherBench 2 framework (Rasp et al., 2020). These notably include the models from the researchers at Google DeepMind (Graphcast, GCM), Nvidia (FourCastNet FCN), and Huawei (Pangu-Weather PGW).
Paper / Scoring Rule Prob. RMSE MAE ACC CRPS SSR Act. Bias RQE Cal. ROC MPE SEEPS FourCastNet (Pathak et al., 2022) GNN (Keisler, 2022) GraphCast (Lam et al., 2023) NeuralGCM (Kochkov et al., 2024) FuXi (Chen et al., 2023) Pangu-Weather (Bi et al., 2023) Pangu-Weather Op. (Bouallègue et al., 2024a) Stormer (Nguyen et al., 2023) a AIFS (Lang et al., 2024a) b Pangu-Weather Op. refers to the operational version of Pangu-Weather. Abbreviations: Prob. (Probabilistic), Act. (Activity), Cal. (Calibration). a Diagnostics not evaluated on the ensemble version of Stormer. b AIFS is in ongoing development for ensemble forecasting.
Table A.1: Summary of included scoring rules across nine significant ML-based weather prediction models (as identified by WeatherBench 2). Abbreviations: ACC (Anomaly Correlation Coefficient), CRPS (Continuous Ranked Probability Score), SSR (Skill Spread Ratio), RQE (Relative Quantile Error), ROC (Receiver Operating Characteristic), MPE (Mean Position Error), and SEEPS (Stable Equitable Error in Probability Space).
@@ A.1 RMSE
with $\theta_{i}^{u}$ and $\theta_{i}^{l}$ being the upper and lower latitude bounds for the grid cell with latitude index (center) $i$. Polar regions, which would require a different formula, are excluded from the 64 by 32 5.625 resolution ERA5 dataset, which is used for training.
For use in evaluation to standardise RMSE results across variables, we introduce the normalised root mean square error NRMSE (Pacchiardi et al., 2024):
@@ A.2 $R^{2}$
where $\bar{o}^{v}$ represents the observation mean. $R^{2}$ in this formulation is bounded above by 1, when forecasts equal observations across all times, $f^{v}_{t,i,j}=o^{v}_{t,i,j}$, but is unbounded from below, due to the absence of an intercept in the neural network. $R^{2}=0$ suggests a deterministic model which does no better than predicting the mean for every observation, with $R^{2}<0$, therefore suggesting poor model fit. We include this metric in evaluation to complement RMSE in assessing accuracy of the forecast mean predictions.
@@ A.3 CRPS
We next consider standard probabilistic scoring rules, designed for ensemble forecasts. The continuous ranked probability score (CRPS), which is strictly proper, is particularly relevant for the probabilistic forecasting of rainfall, due to rainfall having a positive point mass at 0. While other scoring rules for continuous variables are restricted to predictive densities, CRPS instead uses predictive cumulative distributions (Gneiting and Raftery, 2007). For deterministic forecasts, the CRPS reduces to MAE. The standard formulation for CRPS for a variable $v$, with the cumulative distribution function of its probabilistic forecast F is as follows:
@@
We evaluate the NCRPS across specific positions, and normalise by the standard deviation of observations. This normalisation is used to assess the quality of the forecast across the scale of the data, along with adding greater interpretation to CRPS results. A NCRPS greater than 1 represents a forecast error larger than the typical variation, standard deviation, of the observations. NCRPS smaller than 1 means the forecast error is smaller than the typical variation, suggesting reasonable forecasting performance, especially as NCRPS decreases to zero. This relates the performance to the historical spread of climatology in the region, which will be roughly equal to 1.